# Conformal Prediction with Upper and Lower Bound Models

### v2

@@ 3.2 The CPUL approach
The proposed CPUL framework combines several steps to fully exploit the knowledge that $\hat{B}^{l}$ and $\hat{B}^{u}$ provide valid bounds on the target variable, and to account for the heteroskedasticity of their residuals. The algorithm is presented in the NCP framework and, for ease of reading, the presentation omits the strengthening of prediction intervals described in Eq. (14), i.e., strengthening is always performed implicitly. Recall that, by Theorem 3.1, this does not affect the conformalization procedure.
@@
The result follows directly from Theorem 1 from yang2024selection. ∎
@@ 3.3 Relation to other methods
The $\hat{C}^{\text{ul}}$ construction is also related to the adaptive approach followed in sesia2023conformal. The main difference between the two is the use of empirical quantiles $\hat{Q}^{l},\hat{Q}^{u}$ in CPUL, which can be obtained much more efficiently, compared to training additional quantile regression models as in sesia2023conformal. The latter approach provides finer local adaptivity, albeit at a higher computational cost.
@@
$\hat{C}^{**}:=\underset{\hat{C}\in\{\hat{C}^{\text{ll}}_{\tau^{\text{ll}}},\hat{C}^{\text{lu}}_{\tau^{\text{lu}}},\hat{C}^{\text{ul}}_{\tau^{\text{ul}}},\hat{C}^{\text{uu}}_{\tau^{\text{uu}}}\}}{\text{argmin}}\frac{1}{|\mathcal{D}_{\text{cal}}|}\sum_{X_{i}\in\mathcal{D}_{\text{cal}}}\left|\hat{C}(X_{i})\right|$
@@ 4 Optimal Minimal Length Threshold
A key limitation of CPUL, like most well-known CP methods, is its tendency to produce overly narrow prediction intervals, particularly when the initial bounds $\hat{B}^{u}(x)$ and $\hat{B}^{l}(x)$ are tight. The length of the prediction intervals in Eq. (18) scales with the magnitude of $\hat{B}^{u}(x)-\hat{B}^{l}(x)$. Consequently, when the bounds are extremely tight, the constant adjustment of $t$ in the NCP framework becomes disproportionately large, leading empty prediction intervals and, in turn, under-coverage in such regions. Figure 1(b) illustrates this issue in the CPUL-lu setting. After calibrating the initial bounds the area around $x_{0}$ remains uncovered due to the empty intersection of the calibrated intervals. This behavior results in inefficient prediction intervals, as the method should over-cover the less confident regions (i.e., areas where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is large). Hence, the regions with tightest initial bounds become most vulnerable and become under-covered. sesia2020comparison proposed a variant, CQR-r, with $\hat{C}_{t}^{\text{CQR-r}}(x)=\left[\hat{L}_{t}^{\text{CQR-r}}(x),\hat{U}_{t}^{\text{CQR-r}}(x)\right],\ \text{for }x\in\mathcal{X},$ where
@@
The design of sesia2020comparison scales $t$ by $\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)$, which mitigates the disproportionate reduction in prediction interval size. However, as noted in sesia2020comparison, its overall efficiency is worse than the original fixed-length version (12). Thus, it remains unclear whether such scaling, despite being adaptive to the initial tightness of the base models, is preferred over the fixed-length adjustment $t$. Section 5 includes a CQR-r adaptation tailored to the setting of this paper for comparison.
To address this paradoxical miscoverage, the paper proposes the optimal minimal length threshold method (OMLT) as an alternative to the relative scaling approach used in sesia2020comparison. The core idea of OMLT consists in introducing a threshold $\ell\geq 0$, representing the minimum allowed length for a prediction interval. OMLT identifies regions where the given $\hat{B}^{u}$ and $\hat{B}^{l}$ are tightest, where even minor shrinking during calibration poses a high risk of significant undercoverage below the desired level of $1-\alpha$. The design of OMLT retains the standard fixed-length adjustment of $t$ and introduces a threshold for the minimal allowed prediction interval length, effectively marking the boundary of high-risk regions.
@@
The key idea underlying OMLT is that the size of the prediction interval should not be smaller than $\ell$, which prevents under-coverage when a prediction interval is too small. The only exception is for the tightest regions, i.e., when $\hat{B}^{u}(x)-\hat{B}^{l}(x)\leq\ell$, in which case there is no need to enlarge $\bar{C}_{\ell,t}(x)$ beyond $\ell$. The original upper and lower bounds can be confidently relied upon, as $[\hat{B}^{l}(x),\hat{B}^{u}(x)]$ already provides high-quality coverage.
Since (26) reduces to the original problem in (4) when $\ell=0$ for arbitrary $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, the formulation of CPUL-OMLT is designed to produce prediction intervals whose expected length is generally not greater than that of CPUL, at the same $1-\alpha$ coverage. Moreover, OMLT has the potential to extend beyond this setting and reduce the average prediction interval length in cases where disproportionate prediction calibration is observed. For details of implementation, refer to Section A.4.2.
@@ 5 Experiments
The performance of CPUL and CPUL-OMLT are evaluated to demonstrate their ability to achieve valid coverage while producing narrower prediction intervals. The experiments focus on UQ for the optimal value of economic dispatch problems, detailed in Appendix A.1. UQ is conducted given lower and upper bounds derived from primal-dual optimization proxy models chen2023end; qiu2024dual; klamkin2024dual; chen2024real. Namely, the dual proxy provides valid lower bounds ($\hat{B}^{l}$), and the primal proxy provides valid upper bounds $\hat{B}^{u}$). Further details on these optimization proxies are provided in Appendices A.1 and A.3.
@@
Experiments are performed over three datasets: 118_ieee ieee_pstca, 1354_pegase pegase, and 89_pegase pegase, corresponding to power grids of different sizes. For each dataset, samples are randomly shuffled and split 10 times. Dataset details are provided in Section A.2. Evaluation metrics reported are the mean and standard deviation across these 10 runs.
@@
The performance of CPUL-OMLT is compared against several CP baselines. Recall that, unless specified otherwise, all prediction intervals are strengthened as per Eq. (14). The CP baselines include: Split CP (Eq. 6), using either $\hat{B}^{l}$ or $\hat{B}^{u}$ as base predictors; CQR and its variant CQR-r, wherein $\hat{B}^{l}$ and $\hat{B}^{u}$ are treated as the initial lower and upper quantile regressors in the CQR construction; an adapted SFP CP approach of sesia2023conformal. Note that the latter matches the $\hat{C}^{\text{ul}}$ construction. Additional implementation details are provided in Appendix A.4.1.
@@
At a given confidence level $1\,{-}\,\alpha$, shorter interval lengths are considered more desirable. This is captured through expected normalized length of prediction interval $\hat{C}$, defined as $\hat{E}_{X}(\hat{C})=|\mathcal{I}_{\text{test}}|^{-1}\sum_{i\in\mathcal{I}_{\text{test}}}\big{(}\left|Y_{i}\right|^{-1}|\hat{C}(X_{i})|\big{)}$, where $Y_{i}$ and $\hat{C}(X_{i})$ denote the (true) optimal value and the prediction interval for sample $i$. The paper considers the scaled interval length $|\hat{C}(X_{i})|/|Y_{i}|$, expressed as a percentage, rather than absolute interval length $|\hat{C}(X_{i})|$, to account for the possibly large range of values taken by $Y$. This metric is routinely used in the optimization literature chen2023end; vivas2020systematic; wang2009comparison, where it is referred to as optimality gap. All metrics are reported as the mean and standard deviation over 10 runs.
@@
Table 1 and Figure 2 present the numerical performance of the various methods. Table 1 presents, for every dataset, the average and standard deviation of each method’s interval length and coverage, both expressed as a percentage. The table reports results for $\alpha=10\%$; additional $\alpha$ values are reported in Table 3 in Appendix B. Figure 2 displays the interaction between each method’s interval length and coverage, across a broader range of $\alpha$ values. This provides a more global view of each method’s overall performance.
The results in Table 1 demonstrate that all CP methods achieve satisfactory coverage levels, i.e., test coverage is typically close to $1-\alpha$, which is expected given that all methods are properly calibrated. Moreover, all prediction intervals yield a substantial reduction in size compared to the original prediction $[\hat{B}^{l},\hat{B}^{u}]$. For instance, a reduction of only 10% in coverage (i.e., 90% prediction intervals) can yield up to a two- and three-fold reduction in interval size. This demonstrates the value of using UQ techniques to provide more actionable information to practitioners.
@@
Finally, while CPUL and CPUL-OMLT perform similarly on the 89_pegase and 118_IEEE datasets, CPUL-OMLT offers a clear improvement on the 1354_pegase dataset, especially for coverage levels above 90% (see Figure 2). A more granular analysis of each model’s behavior reveals that CPUL’s performace is worse on samples where $[\hat{B}^{l},\hat{B}^{u}]$ is small. Namely, CPUL achieves a coverage of only about 50% across the 5% of test samples with smallest initial interval $[\hat{B}^{l},\hat{B}^{u}]$. Recall that these samples correspond to regions where the given bounds $\hat{B}^{l},\hat{B}^{u}$ are the most accurate, which demonstrates the issue of paradoxical miscoverage (see Section 4 and Figure 1(b)). CPUL-OMLT effectively mitigates this issue, which results in better coverage and smaller interval length overall.

### v3

@@ 3.2 The CPUL approach
The proposed CPUL framework combines several steps to fully exploit the knowledge that $\hat{B}^{l}$ and $\hat{B}^{u}$ provide valid bounds on the target variable, and to account for the heteroskedasticity of their residuals. The algorithm is presented in the NCP framework and, for ease of reading, the presentation omits the strengthening of prediction intervals described in Eq. (14), i.e., strengthening is always performed implicitly. Recall that, by Theorem 1, this does not affect the conformalization procedure.
@@
The result follows directly from Theorem 1 from Yang and Kuchibhotla (2024). ∎
@@ 3.3 Relation to other methods
The $\hat{C}^{\text{ul}}$ construction is also related to the adaptive approach followed in Sesia et al. (2023). The main difference between the two is the use of empirical quantiles $\hat{Q}^{l},\hat{Q}^{u}$ in CPUL, which can be obtained much more efficiently, compared to training additional quantile regression models as in Sesia et al. (2023). The latter approach provides finer local adaptivity, albeit at a higher computational cost.
@@
[…]
@@ 4 Optimal Minimal Length Threshold
A key limitation of CPUL, like most well-known CP methods, is its tendency to produce overly narrow prediction intervals, particularly when the initial bounds $\hat{B}^{u}(x)$ and $\hat{B}^{l}(x)$ are tight. The length of the prediction intervals in Eq. (16) scales with the magnitude of $\hat{B}^{u}(x)-\hat{B}^{l}(x)$. Consequently, when the bounds are extremely tight, the constant adjustment of $t$ in the NCP framework becomes disproportionately large, leading empty prediction intervals and, in turn, under-coverage in such regions. Figure 1(b) illustrates this issue in the CPUL-lu setting. After calibrating the initial bounds the area around $x_{0}$ remains uncovered due to the empty intersection of the calibrated intervals. This behavior results in inefficient prediction intervals, as the method should over-cover the less confident regions (i.e., areas where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is large). Hence, the regions with tightest initial bounds become most vulnerable and become under-covered. Sesia and Candès (2020) proposed a variant, CQR-r, with $\hat{C}_{t}^{\text{CQR-r}}(x)=\left[\hat{L}_{t}^{\text{CQR-r}}(x),\hat{U}_{t}^{\text{CQR-r}}(x)\right],\ \text{for }x\in\mathcal{X},$ where
@@
The design of Sesia and Candès (2020) scales $t$ by $\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)$, which mitigates the disproportionate reduction in prediction interval size. However, as noted in Sesia and Candès (2020), its overall efficiency is worse than the original fixed-length version (12). Thus, it remains unclear whether such scaling, despite being adaptive to the initial tightness of the base models, is preferred over the fixed-length adjustment $t$. Section 5 includes a CQR-r adaptation tailored to the setting of this paper for comparison.
To address this paradoxical miscoverage, the paper proposes the optimal minimal length threshold method (OMLT) as an alternative to the relative scaling approach used in Sesia and Candès (2020). The core idea of OMLT consists in introducing a threshold $\ell\geq 0$, representing the minimum allowed length for a prediction interval. OMLT identifies regions where the given $\hat{B}^{u}$ and $\hat{B}^{l}$ are tightest, where even minor shrinking during calibration poses a high risk of significant undercoverage below the desired level of $1-\alpha$. The design of OMLT retains the standard fixed-length adjustment of $t$ and introduces a threshold for the minimal allowed prediction interval length, effectively marking the boundary of high-risk regions.
@@
The key idea underlying OMLT is that the size of the prediction interval should not be smaller than $\ell$, which prevents under-coverage when a prediction interval is too small. The only exception is for the tightest regions, i.e., when $\hat{B}^{u}(x)-\hat{B}^{l}(x)\leq\ell$, in which case there is no need to enlarge $\bar{C}_{\ell,t}(x)$ beyond $\ell$. The original upper and lower bounds can be confidently relied upon, as $\hat{B}^{l}(x),\hat{B}^{u}(x)]$ already provides high-quality coverage.
Since (24) reduces to the original problem in (4) when $\ell=0$ for arbitrary $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, the formulation of CPUL-OMLT is designed to produce prediction intervals whose expected length is generally not greater than that of CPUL, at the same $1-\alpha$ coverage. Moreover, OMLT has the potential to extend beyond this setting and reduce the average prediction interval length in cases where disproportionate prediction calibration is observed. For details of implementation, refer to Section LABEL:sec:CPUL_and_OMLT_implementation_details.
@@ 5 Experiments
The performance of CPUL and CPUL-OMLT are evaluated to demonstrate their ability to achieve valid coverage while producing narrower prediction intervals. The experiments focus on UQ for the optimal value of economic dispatch problems, detailed in Appendix LABEL:sec:background:optimization. UQ is conducted given lower and upper bounds derived from primal-dual optimization proxy models Chen et al. (2023); Qiu et al. (2024); Klamkin et al. (2024); Chen et al. (2024). Namely, the dual proxy provides valid lower bounds ($\hat{B}^{l}$), and the primal proxy provides valid upper bounds $\hat{B}^{u}$). Further details on these optimization proxies are provided in Appendices LABEL:sec:background:optimization and LABEL:sec:background:proxies.
@@
Experiments are performed over three datasets: 118_ieee University of Washington, Dept. of Electrical Engineering (1999), 1354_pegase Fliscounakis et al. (2013), and 89_pegase Fliscounakis et al. (2013), corresponding to power grids of different sizes. For each dataset, samples are randomly shuffled and split 10 times. Dataset details are provided in Section LABEL:data123. Evaluation metrics reported are the mean and standard deviation across these 10 runs.
@@
The performance of CPUL-OMLT is compared against several CP baselines. Recall that, unless specified otherwise, all prediction intervals are strengthened as per Eq. (14). The CP baselines include: Split CP (Eq. 6), using either $\hat{B}^{l}$ or $\hat{B}^{u}$ as base predictors; CQR and its variant CQR-r, wherein $\hat{B}^{l}$ and $\hat{B}^{u}$ are treated as the initial lower and upper quantile regressors in the CQR construction; an adapted SFP CP approach of Sesia et al. (2023). Note that the latter matches the $\hat{C}^{\text{ul}}$ construction. Additional implementation details are provided in Appendix LABEL:app:experiments:cp_baselines.
@@
At a given confidence level $1\,{-}\,\alpha$, shorter interval lengths are considered more desirable. This is captured through expected normalized length of prediction interval $\hat{C}$, defined as $\hat{E}_{X}(\hat{C})=|\mathcal{I}_{\text{test}}|^{-1}\sum_{i\in\mathcal{I}_{\text{test}}}\big(\left|Y_{i}\right|^{-1}|\hat{C}(X_{i})|\big)$, where $Y_{i}$ and $\hat{C}(X_{i})$ denote the (true) optimal value and the prediction interval for sample $i$. The paper considers the scaled interval length $|\hat{C}(X_{i})|/|Y_{i}|$, expressed as a percentage, rather than absolute interval length $|\hat{C}(X_{i})|$, to account for the possibly large range of values taken by $Y$. This metric is routinely used in the optimization literature Chen et al. (2023); Vivas et al. (2020); Wang et al. (2009), where it is referred to as optimality gap. All metrics are reported as the mean and standard deviation over 10 runs.
@@
Table 1 and Figure 2 present the numerical performance of the various methods. Table 1 presents, for every dataset, the average and standard deviation of each method’s interval length and coverage, both expressed as a percentage. The table reports results for $\alpha=10\%$; additional $\alpha$ values are reported in Table LABEL:table_full:results in Appendix LABEL:sec:Additional_experiment_results. Figure 2 displays the interaction between each method’s interval length and coverage, across a broader range of $\alpha$ values. This provides a more global view of each method’s overall performance.
The results in Table 1 demonstrate that all CP methods achieve satisfactory coverage levels, i.e., test coverage is typically close to $1-\alpha$, which is expected given that all methods are properly calibrated. Moreover, all prediction intervals yield a substantial reduction in size compared to the original prediction $\hat{B}^{l},\hat{B}^{u}]$. For instance, a reduction of only 10% in coverage (i.e., 90% prediction intervals) can yield up to a two- and three-fold reduction in interval size. This demonstrates the value of using UQ techniques to provide more actionable information to practitioners.
@@
Finally, while CPUL and CPUL-OMLT perform similarly on the 89_pegase and 118_IEEE datasets, CPUL-OMLT offers a clear improvement on the 1354_pegase dataset, especially for coverage levels above 90% (see Figure 2). A more granular analysis of each model’s behavior reveals that CPUL’s performace is worse on samples where $\hat{B}^{l},\hat{B}^{u}]$ is small. Namely, CPUL achieves a coverage of only about 50% across the 5% of test samples with smallest initial interval $\hat{B}^{l},\hat{B}^{u}]$. Recall that these samples correspond to regions where the given bounds $\hat{B}^{l},\hat{B}^{u}$ are the most accurate, which demonstrates the issue of paradoxical miscoverage (see Section 4 and Figure 1(b)). CPUL-OMLT effectively mitigates this issue, which results in better coverage and smaller interval length overall.